In 7 years harry will be two times as old as he was last year how old is he now
Answer by addingup(3677)

h+7 = 2(h-1)
h+7 = 2h-2
2h-2 = h+7
add 2 and subtract h on both sides:
h = 9
-------------------
Check:
9+7 = 16
last year Harry was 9-1 = 8 and 2 times Harry last year = 16
Correct!